# ILID: Native Script Language Identification for Indian Languages

### v1

@@ Abstract
Language Identification task is a crucial fundamental step in NLP. Often it serves as a pre-processing step for widely used NLP applications such as multilingual machine translation, information retrieval, question and answering, and text summarization. The core challenge of language identification lies in distinguishing languages in noisy, short, and code-mixed environments. This becomes even harder in case of diverse Indian languages that exhibit lexical and phonetic similarities, but have distinct differences. Many Indian languages share the same script making the task even more challenging. In this paper, we release a dataset of 230K sentences consisting of English and all 22 official Indian languages labeled with their language identifiers where data in most languages are newly created. We also develop and release robust baseline models using state-of-the-art approaches in machine learning and deep learning that can aid the research in this field. Our baseline models are comparable to the state-of-the-art models for the language identification task.
@@ 1 Introduction
India is a linguistically diverse country. Although there are more than 1000 languages in the Indian subcontinent, the digital divide among many Indian languages is enormous. Barring a few languages, most of the languages suffer from resource scarcity and a simple task such as Language IDentification (LID) remains a challenging task Caswell et al. (2020) for them. Available LID tools such as Lui and Baldwin (2011), Joulin et al. (2016a), and Team et al. (2022) perform poorly on Indian languages and they do not cover many of them. With the increase in mobile and internet users in India, the need for providing services in native Indian languages is more pressing than ever. Hence, it becomes essential to create robust LID tools to cater to the Indian users. As the volume and variety of data keeps growing, we need to include data from different domains while creating sound benchmarks. Keeping this motivation in mind, we set out to create an Indian Language IDentification (ILID) benchmark including English and 22 official languages consisting of a total 230K sentences labeled with their language markers. The main contribution of our paper is two folds:
@@
  • We create ILID dataset consisting of newly created datasets in 13 languages and curated datasets in rest of the languages.
@@
  • We develop ILID baseline models exploring different machine learning and deep learning techniques.
@@ 2 ILID Dataset Creation
Indian Language IDentification (ILID) dataset is created using two approaches. We include English and all 22 official Indian languages  [^1] https://en.wikipedia.org/wiki/Eighth_Schedule_to_the_Constitution_of_India widely used in India. The first approach utilizes web scraping for the languages in which the digital presence is significant (details are given in the Appendix). For each of these languages, we collect 10,000 different sentences from diverse sources such as government websites, newspapers, books, and other public materials ensuring varying degrees of linguistic complexity. The second approach samples sentences from an existing massive monolingual and parallel corpora for Indian languages Mujadia and Sharma (2024). The details of the dataset are presented in Table 1. The data in each language is split into 80:10:10 ratio to create train, dev, and test sets.
@@
Manipuri (mni) 8000 1000 1000 10000
@@
Sindhi (snd) 8000 1000 1000 10000
@@
To ensure quality and consistency, the dataset undergoes several noise removal steps. The first step involves the elimination of duplicate, very short, and ungrammatical sentences. The next step employs an existing FastText Joulin et al. (2016b, a) based language identification model to remove sentences where the probability of detecting the language of the text is very low (we fix a threshold of 0.7 for this filtering process). This happens in code-mixed texts as many Indians are multilingual  [^2] https://en.wikipedia.org/wiki/List_of_languages_by_number_of_native_speakers_in_India and often use more than one language while writing.
The first approach is described in details.
@@ 2.1 Web Scraping
To build a strong language identification system for Indian languages, we carefully built a personalized set of 10,000 text samples for each language, covering 12 Indian languages. Details are available in Appendix A Table 4.
@@ 2.2 Data Cleaning Pipeline
In the data cleaning process we mainly work on removing extra spaces, extra special symbols, unicode normalization, breaking the paragraph into meaningful sentences, and after that we apply various tokenization methods. For sentence tokenization, we utilize the corresponding end sentence markers defined for respective languages. We use a regular expression based tokenizer  [^3] https://github.com/Pruthwik/Tokenizer_for_Indian_Languages for sentence and word tokenization. Our cleaning process transforms raw text $T$ to cleaned text $T^{\prime}$ through:
@@
where $\psi_{s}$ handles special characters and $\psi_{n}$ performs normalization. We use a sentence as a unit for the LID task. Each sentence is labeled with a language tag. ISO 639-2 codes  [^4] https://en.wikipedia.org/wiki/List_of_ISO_639-2_codes or 3 lettered language codes denote each language detailed in Appendix B Table 5. In the second approach, sentences are randomly sampled from the huge corpora Mujadia and Sharma (2024) that contain the information about the language. A similar exercise of appending the labels to the sentences as language identifiers and data splitting is performed on these samples.
3 ILID Model
The ILID model is a classifier that can categorize a piece of text into one of the 23 classes that includes English and 22 Indian languages. The architecture of the proposed system is shown in Figure 1. We implement three kinds of approaches for designing the classifiers.
3.1 Machine Learning Models
In this approach, each sentence is represented by a TF-IDF Sparck Jones (1972) vector in a bag-of-words setting. We utilize both word-level and character-level TF-IDF representations to enhance textual feature extraction. Word-level TF-IDF captures the importance of entire words across the corpus, which helps in preserving semantic meaning and understanding context at a higher level. However, word-level models may struggle with noisy or morphologically rich languages, spelling errors, or out-of-vocabulary terms. To address this, we also incorporate character-level TF-IDF, which analyzes smaller units such as character n-grams. This enables the model to recognize sub-word patterns and partial word structures, making it more resilient to variations and linguistic nuances. By combining both approaches, our system is able to capture both the global meaning (through words) and the finer-grained patterns (through characters), resulting in a more robust and language-agnostic representation, especially beneficial for multilingual or noisy text data. Eight different classifiers such as Support Vector Machine (SVM), Logistic Regression (LR), Random Forest (RF), Decision Tree (DT), Naive Bayes’, Stochastic Gradient Descent (SGD), K-Nearest Neighbors, and Adaboost are explored for this task. We develop different ensemble models consisting of 3, 4, and 5 classifiers and evaluate their performance. Due to paucity of space, we only report the five best performing ensemble models.
Model KNN DT RF SVM NB LogReg SGD F1 (Dev) F1 (Test)
Voting-1 0.99 0.96
Voting-2 0.99 0.96
Voting-3 0.99 0.96
Voting-4 0.99 0.96
Voting-5 0.99 0.99
MuRIL 0.96 0.96
FastText 0.96 0.96
Table 2: Comparison of Voting Classifiers, MuRIL, and FastText on Dev and Test sets
Model KNN DT RF SVM NB LogReg SGD ILID IndicLID
Voting-1 0.98 0.94
Voting-2 0.98 0.96
Voting-3 0.98 0.95
Voting-4 0.98 0.95
Voting-5 0.98 0.97
MuRIL 0.98 0.94
FastText 0.98 0.92
Table 3: Comparison of Voting Classifiers, MuRIL, and FastText on Other Benchmarks
3.2 FastText Classifier
FastText Joulin et al. (2016b, a) utilizes word embeddings Bojanowski et al. (2017) composed of character n-grams that can better represent rare words and orthographic similarities. FastText classifier is a linear classifier and is very fast at computation.
4 Pretrained BERT Model
Pretrained subword based contextual language models such as BERT Devlin et al. (2019), XLM Conneau et al. (2020), RoBERTa Zhuang et al. (2021) have proven to be very effective in text classification tasks or generally natural language understanding tasks. One variant of BERT, MuRIL Khanuja et al. (2021) is pretrained on Indian languages data. Hence, we fine-tune the MuRIL pretrained model on the ILID train set for our task.
5 Experimental Details
The machine learning models have been implemented using the Scikit-learn Pedregosa et al. (2011) framework. Similarly, the FastText library  [^5] https://fasttext.cc/docs/en/supervised-tutorial.html is utilized to implement the classification on the languages of the texts. The MuRIL models is fine-tuned using the Hugginggace transformer Wolf et al. (2019) framework. We will upload the models with MIT license once the paper gets accepted.
6 Evaluation Metrics
All models have been evaluated using the macro F1 scores. The macro F1 score averages the F1 scores across all the languages. In order to evaluate the score, the precision and recall scores are also computed at the macro level. Each ensemble employs majority voting mechanism to determine the predicted labels.
7 Results and Discussion
The ensemble machine learning models perform better than the individual models. The performance of the ensembles is also superior to the FastText and fine-tuned MuRIL models. The best five performing ensembles along with the FastText and fine-tuned MuRIL models are presented in Table 2 that represent the macro F1 score for each model. The results for the individual models are shown in Tables 6 and 7. Although the scores are encouraging, the performance drops in languages that share a script such as Devanagari for Hindi, Maithili, Marathi, Konkani, and Sanskrit, Arabic for Kashmiri, Urdu, and Sindhi. We compare our models with the state-of-the-art IndicLID model Madhani et al. (2023) on the Bhasha-Abhijnaanam Madhani et al. (2023) benchmark where our model perform comparatively. Our ILID model perform better on the languages where data are scraped while the performance dips for languages that are created from other external resources.
8 Conclusion
In this paper, we create the ILID benchmark dataset for English and Indian languages and develop baseline language identification models. The dataset and models will serve as foundation tools mostly for the resource poor languages that are part of this benchmark. We hope that this will act as an impetus for future NLP research in these languages.
Appendix A Languages Used in Scraping and Sampling
As explained in 2, based on their online availability, we used two approaches of scraping and sampling from an existing corpus for the dataset creation. The languages are shown in Table 4.
@@ Appendix A Languages Used in Scraping and Sampling
Appendix B ISO 639-2 Codes
We utilize ISO 639-2 codes as the language labels for this task.
@@ Appendix B ISO 639-2 Codes
Manipuri mni
@@
Sindhi snd

### v2

@@ Abstract
The language identification task is a crucial fundamental step in NLP. Often it serves as a pre-processing step for widely used NLP applications such as multilingual machine translation, information retrieval, question and answering, and text summarization. The core challenge of language identification lies in distinguishing languages in noisy, short, and code-mixed environments. This becomes even harder in case of diverse Indian languages that exhibit lexical and phonetic similarities, but have distinct differences. Many Indian languages share the same script, making the task even more challenging. Taking all these challenges into account, we develop and release a dataset of 250K sentences consisting of 23 languages including English and all 22 official Indian languages labeled with their language identifiers, where data in most languages are newly created. We also develop and release baseline models using state-of-the-art approaches in machine learning and fine-tuning pre-trained transformer models. Our models outperforms the state-of-the-art pre-trained transformer models for the language identification task. The dataset and the codes are available at https://yashingle-ai.github.io/ILID/ and in Huggingface open source libraries.
@@ 1 Introduction
India is a linguistically diverse country. Although there are more than 1000 languages in the Indian subcontinent, the digital divide among many Indian languages is enormous. Barring a few languages, most of the languages suffer from resource scarcity and a simple task such as Language IDentification (LID) remains a challenging task Caswell et al. (2020) for them. Available LID tools such as Lui and Baldwin (2011), Joulin et al. (2016a), and Team et al. (2022) perform poorly on Indian languages and they do not cover many of them. With the increase in mobile and internet users in India, the need for providing services in native Indian languages is more pressing than ever. Hence, it becomes essential to create robust LID tools to cater to Indian users. As the volume and variety of data continue to grow, we need to include data from different domains while creating sound benchmarks. Keeping this motivation in mind, we set out to create an Indian Language Identification (ILID) benchmark including English and 22 official languages spanning 25 scripts consisting of a total of 250K sentences labeled with their language markers. The main contribution of our paper is two fold.
@@
  • We create the ILID dataset consisting of newly created datasets in 13 languages and curated datasets in the remaining ten languages.
@@
  • We develop ILID baseline models using different machine learning and deep learning techniques.
  • We perform a comparative analysis of our models with the state-of-the-art model.
2 Related Work
Language Identification (LID) is a fundamental task in Natural Language Processing (NLP), aiming to determine the language of a given text or speech segment. Early approaches to LID primarily relied on statistical methods, most notably character $n$-gram models Cavnar and Trenkle (1994). These methods leverage the unique frequency distributions of character sequences within different languages to classify unknown texts. Language Identification becomes more crucial in the domain of Cross Lingual Information Retrieval where the query is written in native scripts Bosca and Dini (2010). Similarly, spoken language identificationDehak et al. (2011); Gonzalez-Dominguez et al. (2014) is a very essential task in multilingual speech processing that can further impact in developing applications such as machine translation and speech recognition. In this paper, we are dealing with the language identification task in written texts. With machine learning approaches such as Support Vector Machines (SVMs) Dunning (1994), Naive Bayes classifiers Elworthy (1998), decision trees, random forests, and gradient boosting gaining widespread usage, LID systems were also modeled to leverage these techniques with the same $n$-gram features and other lexical or language specific hand crafted features. While effective for well-resourced languages with distinct orthographies, the performance of the statistical systems can degrade when dealing with short texts, noisy data, or languages sharing common scripts. Deep learning methods such as Convolutional Neural Networks (CNNs) Kim (2014) and Recurrent Neural Networks (RNNs) Hochreiter and Schmidhuber (1997) (particularly LSTMs and GRUs) are often trained on distributionally similar word or character n-gram or single character embeddings vectors to better capture intricate language complexity and handle out-of-vocabulary words more robustly Conneau et al. (2017). The current approaches include fine-tuning pre-trained transformer Vaswani et al. (2017) models using subwords Sennrich et al. (2016); Kudo and Richardson (2018); Song et al. (2021). Many of the language identification tasks are modeled as token classification tasks that are helpful in code-mixed settings.
The task of language identification for Indian languages is even more challenging as they belong to different and diverse language families (Indo-Aryan, Dravidian, Tibeto-Burman, Austroasiatic), and written in over a dozen distinct scripts. Multiple Indian languages often share the same script. For example; Hindi, Marathi, Nepali, and Sanskrit are all primarily written in the Devanagari script. This script overlap requires models to distinguish languages based on lexical, morphological, or syntactic features rather than solely on orthographic cues.
Another prevalent issue in Indian language contexts is code-mixing and code-switching, where speakers or writers regularly use words or phrases from multiple languages within a single utterance or text Gambäck and Das (2017). This phenomenon is common in informal communication and social media Barman et al. (2014), making it difficult for traditional LID systems to accurately identify the primary language or even segment code-mixed segments. Research in this area often focuses on fine-grained LID, aiming to identify language at the word or phrase level Bhat et al. (2014), rather than just document or sentence level identification.
The language identification task for Indian languages also followed the similar line of research using rule-based and machine learning based approaches. More advanced approaches fine-tune pre-train transformers trained on Indian languages Kumar et al. (2023); Agarwal et al. (2023); Madhani et al. (2023a). Madhani et al. (2023a) also transliterate Madhani et al. (2023b) the native scripts into roman scripts for uniformity across languages and better efficiency in the LID systems. Despite these advancements, robust and fine-grained LID for all Indian languages, especially in code-mixed scenarios and for low-resource languages, remains an active area of research. In this paper, we limit the scope of identifying the language at a sentence level, not at the constituent token level.
@@ 3 ILID Dataset Creation
Indian Language Identification (ILID) dataset is created using two approaches. We include English and the 22 official Indian languages  [^1] https://en.wikipedia.org/wiki/Eighth_Schedule_to_the_Constitution_of_India [^2] https://en.wikipedia.org/wiki/Linguistic_Survey_of_India widely used in India. The first approach utilizes web scraping for the languages in which the digital presence is significant (details are given in the Appendix). For each of these languages, we collect 10,000 different sentences from various sources such as government websites, newspapers, books, and other public materials, ensuring varying degrees of linguistic complexity. The second approach samples sentences from Bhashaverse Mujadia and Sharma (2024), an existing massive monolingual and parallel corpora for Indian languages. The details of the dataset are presented in Table 1. The data in each language is split into 80:10:10 ratio to create train, dev, and test sets. The dataset is available at https://huggingface.co/datasets/yash-ingle/ILID_Indian_Language_Identification_Dataset.
@@
Manipuri Bengali Script (mni_Beng) 8000 1000 1000 10000
Manipuri Meitei Script (mni_Mtei) 8000 1000 1000 10000
@@
Sindhi Devnagari(snd_Deva) 8000 1000 1000 10000
Sindhi Perso-Arabic(snd_Arab) 8000 1000 1000 10000
@@
To ensure quality and consistency, the dataset undergoes several noise removal steps. The first step involves the elimination of duplicate, very short, and ungrammatical sentences. The next step employs an existing FastText Joulin et al. (2016b, a) based language identification model to remove sentences where the probability of detecting the language of the text is very low (we fix a threshold of 0.7 for this filtering process). This happens in code-mixed texts as many Indians are multilingual and often use more than one language while writing.
The approaches are described in detail.
@@ 3.1 Web Scraping
To build a strong language identification system for Indian languages, we carefully built a personalized set of 10,000 text samples for each language, covering 13 Indian languages.
@@ 3.2 Data Cleaning Pipeline
In the data cleaning process, we mainly work on removing extra spaces, extra special symbols, unicode normalization, breaking the paragraph into meaningful sentences, and after that we apply various tokenization methods. For sentence tokenization, we utilize the corresponding end sentence markers defined for respective languages. We use a regular expression based tokenizer  [^3] https://github.com/Pruthwik/Tokenizer_for_Indian_Languages for sentence and word tokenization that is specifically designed to handle texts in Indian languages. Our cleaning process transforms raw text $T$ to cleaned text $T^{\prime}$ through:
@@
where $\psi_{s}$ handles special characters and $\psi_{n}$ performs normalization. We use a sentence as a unit for the LID task. Each sentence is labeled with a language tag. ISO 639-2 codes  [^4] https://en.wikipedia.org/wiki/List_of_ISO_639-2_codes or 3 lettered language codes denote each language detailed in Table 3.
3.3 Sampling From Existing Corpora
In the second approach, sentences are randomly sampled from Bhashaverse Mujadia and Sharma (2024) that contain a huge collection of monolingual corpora of different Indian languages. In this corpora, Manipuri, and Sindhi have sentences in two different scripts each. Manipuri is available in Bengali and Meitei scripts, whereas Sindhi is written in Devanagari and Perso-Arabic scripts. We have sampled sentences using all available scripts from these two languages. This increases the total number of languages with different scripts to 25 also presented in Table 4. A similar exercise of appending the labels to the sentences as language identifiers and data splitting is performed on these samples. The language specifics using both approaches are shown in Table 2.
@@ 3.3 Sampling From Existing Corpora
Manipuri Bengali Script mni_Beng
Manipuri Meitei Scrip mni_Mtei
@@
Sindhi Perso-Arabic snd_Arab
Sindhi Devanagari snd_Deva